# MAMA-MIA: A Large-Scale Multi-Center Breast Cancer DCE-MRI Benchmark Dataset with Expert Segmentations

### v2

@@ Methods
An additional contribution of this work is the pretrained weights of a vanilla nnUNet [14] tumor segmentation model, trained using the 1506 expert segmentations. The nnUNet model was trained with DCE-MRI full-images as input over a total of 1000 epochs in a five-fold cross-validation setting. The model achieved a mean validation Dice coefficient of $0.7620\pm 0.2113$.
@@
All data records, including the DCE-MRI images, the automatic and expert segmentations for each of the 1506 cases in the MAMA-MIA dataset and the weights of the pretrained segmentation model, are available online in the MAMA-MIA Synapse repository (https://doi.org/10.7303/syn60868042). Data records also include three tables, one that contains all the clinical and imaging information, another with the automatic segmentation quality scores from two experts who evaluated the automatic segmentations using the GUI, and a table with the train and test split to promote reproducibility in future studies using the dataset. Figure 1 illustrates the file content and folder structure of the MAMA-MIA dataset. Each case identifier consists of the original collection/dataset acronym and the corresponding patient identification number (patient ID). For instance, the ISPY1_1221 case corresponds to the pre-treatment DCE-MRI sequences of patient ID 1221 from the ISPY-1 collection. The different phases are named using the same case ID plus the corresponding phase number (ISPY1_1221_ [^0] X). For example, ISPY1_1221_0000 represents the pre-contrast phase and ISPY1_1221_0002 represents the second post-contrast phase.
@@
Clinical information available in more than 90% of the cases includes the presence of bilateral cancer at diagnosis, multi-focal cancer, tumor subtype, and pathological complete response (pCR) after neoadjuvant chemotherapy (NAC) treatment. Other relevant information included in the dataset, albeit not present in all cases, comprises survival status in over 450 cases, different tumor receptors, days to recurrence or metastasis, agents prescribed during NAC, the necessity of mastectomy after treatment, and more. A comprehensive list of clinical and imaging variables included in the dataset can be found in the Excel Table as part of the Supplementary Material.
Table 2 presents the most common imaging information included in the dataset: date of original collections, acquisition plane, magnetic field strength (Tesla) used for DCE-MRI acquisition, scanner manufacturers and models, number of bilateral and fat-suppressed DCE-MRIs, mean number of slices, slice thickness, pixel spacing, number of phases, and total number of cases obtained from each original collection. In the DCE-MRI medical imaging modality, the acquisition time interval between contrast administration (pre-contrast MRI) and subsequent post-contrast phases is an important factor. Table 3 summarizes the average time intervals between phases per dataset. It is notable that older datasets like ISPY1 and NACT have longer acquisition intervals than later collections.
Figure 6: From left to right: a) distribution of expert-assigned quality scores for automatic segmentations, b) Dice Similarity Coefficient (DSC) and c) the 95 percentile Hausdorff Distance (HD) between automatic and manual segmentations across different quality categories. The middle panel displays DSC (higher values indicate better agreement), while the right panel shows HD (lower values indicate better alignment), respectively.
@@
Figure 7: Examples of automatic (yellow) and expert segmentations (red) and the corresponding expert corrections (pink if the voxels were removed, and green if they were added). The examples are grouped by the expert-assigned quality scores for automatic segmentations during the visual inspection (a) Missed, b) Poor, c) Acceptable, and d) Good). Under the automatic segmentation there is the Dice Similarity Coefficient (DSC) and the 95 percentile Hausdorff Distance (HD) between automatic and expert segmentations. Also, under the manual segmentation, the final volume of the expert segmentation in $cm^{3}$. The total volume corrected by the experts is quantified in $cm^{3}$.
In this section, we provide a detailed validation of the segmentation quality in our dataset. Expert segmentations are inherently validated through expert review, but we also use preliminary automatic segmentations as a baseline for further validation. To assess the quality of these automatic segmentations, two external expert radiologists evaluated them, categorizing each segmentation into four quality levels: Missed, Poor, Acceptable, and Good. We present a comprehensive analysis of these quality assessments, using distance-based metrics such as the Dice Similarity Coefficient (DSC) and the 95 percentile Hausdorff Distance (HD), to compare the automatic segmentations with the final expert segmentations. Additionally, we explore the implications of these assessments and suggest potential improvements for future segmentation quality control.
A comprehensive list of expert quality scores and corresponding distance metrics is included in the Data Records as a CSV file, facilitating exhaustive analysis of automatic segmentation quality and expert agreement. Figure 6 illustrates the automatic segmentation distributions across different quality scores between the two experts, alongside with the DSC and HD between the final expert segmentations and the preliminary automatic segmentations.
Overall, Expert 1 deemed 669 (44.4%) automatic segmentations as Good quality, and Expert 2 a total of 652 (43.3%) (Figure 6 (a)). Good automatic segmentations correlated with notably high DSC values and minimal HD, suggesting robust agreement with manual segmentations. The efficiency and adequacy of GUI inspection for Good automatic segmentations were validated by these metrics. Automatic segmentations rated as Acceptable showed similar DSC and HD ranges across both experts, closely resembling manual segmentations (Figure 6 (b) and (c)). Expert 1 labeled 192 (12.7%) as Acceptable, while Expert 2 labeled 376 (24.9%). For automatic segmentations categorized as Poor, Expert 1 was more stringent than Expert 2, labeling as Poor a total of 325 (21.5%) of automatic segmentations compared to Expert 2’s 151 (10%). However, the standard deviation of DSC for Poor segmentations by Expert 1 exceeded that of Expert 2, at times surpassing the DSC range of Acceptable segmentations by the same expert. Both experts, considered that Missed tumors are less than 2% of the automatic segmentations, exhibiting lower DSC values and larger HD, indicating large differences from manual segmentations and confirming the bad quality of these segmentations. The disparity in the number of automatic segmentations categorized as Missed between experts suggests that additional information beyond the displayed GUI images may have been necessary to confirm tumor omission. The primary discrepancy between experts in evaluating automatic segmentation quality stemmed from categorization as Poor or Acceptable. Future revisions of segmentation quality control may benefit from reducing categories to three: Missed, Corrections Needed, and Good. This approach could minimize intra-observer variability while identifying automatic segmentations requiring manual correction. We consider the quality scores provided by the two experts for the automatic segmentations to be a valuable resource for studies on automatic segmentation quality control.
Additionally, to illustrate the corrections made by experts to the automatic segmentation, we grouped some examples in Figure 7 based on quality scores. The expert additions to the preliminary automatic segmentations are shown in green, while the voxels removed are depicted in pink. We can observe that for automatic segmentations assessed as Good in the visual quality control, the corrections are minimal, mainly focusing on refining cancer margins and other fine-grain details. Apart from using the Dice and the 95 percentile Hausdorff Distance (HD) metrics to evaluate the quality of the automatic segmentations and the extent of corrections needed, it is interesting to visualize Missed examples like ISPY2_566011, where the main lesion was not segmented, and the expert had to segment the missed lesion from scratch.
Usage Notes
The dataset and data records are hosted in Synapse (https://doi.org/10.7303/syn60868042). In addition, to facilitate the use of this dataset and to get familiarized with the data structure, we have released a Github repository. The repository contains Jupyter notebooks to read the images and visualize them and code to run the inference with the pretrained nnUNet model.
@@
The GitHub repository of this dataset is available in: https://github.com/LidiaGarrucho/MAMA-MIA. The automatic segmentation models were trained using the code from nnUNet GitHub repository in a 5-fold cross validation setting. The original DICOM images were transformed to NifTI using the pycad Python library. The Dice and 95% Hausdorff Distance metrics are computed using seg-metrics 1.2.7 [34] Python library.
@@ Acknowledgements
Margnetic field strength 1.5 171 (100.0) 715 (73.0) 136 (46.7) 64 (100.0) 1086 (72.1)

### v3

@@ Methods
An additional contribution of this work is the pretrained weights of a vanilla nnU-Net [14] tumor segmentation model, trained using the 1506 expert segmentations. The nnU-Net model was trained with DCE-MRI full-images as input over a total of 1000 epochs in a five-fold cross-validation setting. The model achieved a mean validation Dice coefficient of $0.7620\pm 0.2113$.
@@
All data records, including the DCE-MRI images, the automatic and expert segmentations for each of the 1506 cases in the MAMA-MIA dataset and the weights of the pretrained segmentation model, are available online in the MAMA-MIA Synapse repository [36] under CC-BY-NC license at the level of the most restrictive primary dataset (Duke-Breast-Cancer-MRI [28]). Data records also include three tables, one that contains all the clinical and imaging information, another with the automatic segmentation quality scores from two experts who evaluated the automatic segmentations using the GUI, and a table with the train and test split to promote reproducibility in future studies using the dataset. Figure 1 illustrates the file content and folder structure of the MAMA-MIA dataset. Each case identifier consists of the original collection/dataset acronym and the corresponding patient identification number (patient ID). For instance, the ISPY1_1221 case corresponds to the pre-treatment DCE-MRI sequences of patient ID 1221 from the ISPY-1 collection. The different phases are named using the same case ID plus the corresponding phase number (ISPY1_1221_ [^0] X). For example, ISPY1_1221_0000 represents the pre-contrast phase and ISPY1_1221_0002 represents the second post-contrast phase.
@@
Clinical information available in more than 90% of the cases includes the presence of bilateral cancer at diagnosis, multifocal cancer, tumor subtype, and pathological complete response (pCR) after neoadjuvant chemotherapy (NAC) treatment. Other relevant information included in the dataset, albeit not present in all cases, comprises survival status in over 450 cases, different tumor receptors, days to recurrence or metastasis, agents prescribed during NAC, the necessity of mastectomy after treatment, and more. A comprehensive list of clinical and imaging variables included in the dataset can be found in the Excel Table as part of the Supplementary Material.
Figure 6: Examples of preliminary automatic (yellow) and expert segmentations (red) and the corresponding expert corrections (pink if the voxels were removed, and green if they were added). The examples are grouped by the expert-assigned quality scores for automatic segmentations during the visual inspection (a) Missed, b) Poor, c) Acceptable, and d) Good). Under the automatic segmentation there is the Dice Similarity Coefficient (DSC) and the 95 percentile Hausdorff Distance (HD) between automatic and expert segmentations. Also, under the manual segmentation, the final volume of the expert segmentation in $cm^{3}$. The total volume corrected by the experts is quantified in $cm^{3}$.
Table 2 presents the most common imaging information included in the dataset: date of original collections, acquisition plane, magnetic field strength (Tesla) used for DCE-MRI acquisition, scanner manufacturers and models, number of bilateral and fat-suppressed DCE-MRIs, mean number of slices, slice thickness, pixel spacing, number of phases, and total number of cases obtained from each original collection. In the DCE-MRI medical imaging modality, the acquisition time interval between contrast administration (pre-contrast MRI) and subsequent post-contrast phases is an important factor. Table 3 summarizes the average time intervals between phases per dataset. It should be noted that older datasets like ISPY1 and NACT have longer acquisition intervals than later collections.
@@
This technical validation evaluates the visual quality control of preliminary automatic segmentations, emphasizing the role of expert-corrected segmentations as the primary contribution of this study. The goal is to provide actionable guidelines and recommendations for future segmentation quality assessments by analyzing inter-rater agreement between two expert radiologists and correlating their evaluations with established quality metrics, such as the Dice Similarity Coefficient (DSC) and Hausdorff Distance (HD).
Representative examples of expert corrections are shown in Figure 6, categorized by expert-assigned quality scores of the preliminary automatic segmentations. It is important to note that all cases, including those rated as Good during the visual assessment, were sent to the radiologists for manual correction without revealing the quality category (Good, Acceptable, Poor, or Missed). This blinded approach ensured that no bias was introduced during the correction process. Green voxels highlight areas added by the experts, while pink voxels indicate areas removed. Minimal corrections were necessary for segmentations rated as Good, which typically involved refining tumor boundaries. In contrast, cases categorized as Missed, such as ISPY2_566011, required the experts to perform full tumor segmentation from scratch. These examples underscore the critical role of robust automatic segmentation methods in minimizing the burden of manual corrections and highlight the value of expert-corrected segmentations.
Figure 7 illustrates the distribution of quality scores assigned by the experts, alongside the DSC and HD values calculated between expert-corrected and preliminary automatic segmentations. Segmentations rated as Good were strongly correlated with high DSC values and low HD, confirming the reliability of the visual quality control interface in distinguishing segmentations that meet clinical standards. However, the greatest discrepancies between experts occurred in the Acceptable and Poor categories, where subjective interpretation played a larger role in the quality assessment.
Figure 7: From left to right: a) distribution of expert-assigned quality scores for automatic segmentations, b) Dice Similarity Coefficient (DSC) and c) the 95 percentile Hausdorff Distance (HD) between automatic and manual segmentations across different quality categories. Panel (b) displays DSC (higher values indicate better agreement), while Panel (c) shows HD (lower values indicate better alignment), respectively.
To quantify inter-rater agreement, we computed Cohen’s Kappa under two categorization schemes: the original four categories (Good, Acceptable, Poor, and Missed) and a simplified binary scheme (Good versus Corrections Needed). The four-category scheme yielded a Cohen’s Kappa of 0.39, reflecting low to moderate agreement and indicating inconsistencies, particularly in intermediate categories. Simplifying to a binary scheme improved agreement significantly, with a Cohen’s Kappa of 0.53, demonstrating that a simplified approach improves consistency between experts while maintaining the primary goal of identifying segmentations that require manual correction.
These findings emphasize the efficiency of adopting a binary categorization scheme in visual quality control workflows. By focusing only on segmentations requiring correction (Corrections Needed) versus those considered satisfactory (Good), future visual assessments can achieve greater inter-rater agreement, reduce annotation time, and optimize the use of expert resources. The results of this analysis aim to contribute to the development of robust visual quality control guidelines for assessing automatic segmentation models.
Furthermore, the categorization of the preliminary automatic segmentations by the radiologists provides a valuable dataset that can be leveraged to train a classifier capable of mimicking radiologists in assessing segmentation quality. Such a classifier could automatically identify Good segmentations or flag those requiring corrections, further reducing the reliance on manual evaluations and improving scalability. A comprehensive list of expert quality scores and corresponding distance metrics is included in the Data Records as a CSV file, enabling further analysis, benchmarking efforts, and the potential training of automated quality assessment tools.
@@
The GitHub repository for the MAMA-MIA dataset is available at: https://github.com/LidiaGarrucho/MAMA-MIA. This repository provides detailed resources and instructions to support researchers in integrating the MAMA-MIA dataset into their studies. It includes comprehensive scripts to facilitate working with the dataset and enable customization of workflows for segmentation, detection, and analysis in various radiomic and AI-driven studies. The key components of the repository are as follows:
  • Data Filtering and Visualization Scripts: Jupyter notebooks are provided to guide users in filtering and visualizing clinical and imaging data using the pandas library, along with the clinical and imaging variables included in the dataset’s supplementary Excel table. The notebooks include:
    • Filtering patients based on clinical outcomes, such as tumor subtype, treatment response, or survival status.
    • Sorting and grouping cases by scanner manufacturer, image resolution, or pixel spacing.
    • Exploring relevant clinical and imaging metadata from the dataset information table.
    • Displaying filtered datasets in tabular or graphical formats for further analysis.
    • Plotting MRI images using matplotlib for quick visual inspection.
    • Displaying VOIs and segmentation contours, along with 3D visualizations of tumor segmentations.
  • Preprocessing Script: A Jupyter notebook describing common preprocessing steps has been designed to ensure the dataset’s compatibility with a wide range of analysis pipelines, while offering flexibility for researchers to adapt workflows to their specific objectives. For a detailed guide, refer to the notebooks directory. The script includes methods to perform:
    • Bias Field Correction: Corrects for intensity inhomogeneities and biases, which can enhance the accuracy of subsequent analyses.
    • Denoising: A non-linear filter that averages similar patches throughout the image, allowing it to selectively reduce noise without sacrificing edges and subtle intensity variations.
    • Intensity Clipping: clipping intensity values to a defined percentile range can help reduce outliers.
    • Normalization (Z-score): Pixel intensity normalization is performed by subtracting the mean and dividing by the standard deviation, resulting in a distribution with zero mean and unit standard deviation.
    • Histogram Equalization: Enhances the contrast of the images by redistributing the intensity values to span the full range of the histogram.
    • Resampling: e.g. Resampling the images to isotropic voxel size (e.g., $1\times 1\times 1$ mm³), ensuring uniform resolution across all images.
  • Bounding Box Extraction: Code to extract 3D bounding box coordinates containing primary lesions to ease the detection and isolation of VOIs for tumor analysis.
  • nnU-Net Segmentation Model Training: Detailed steps for training an nnU-Net model using the dataset’s segmentations. The automatic segmentation models were trained using the nnU-Net framework in a 5-fold cross-validation setting. Additionally, the pretrained weights of the segmentation model, trained on all expert segmentations, are shared and hosted in the Synapse repository. This facilitates model development and enables easier benchmarking and comparison by other researchers.
The original DICOM images from TCIA were transformed to NIfTI format using the pycad Python library. Metrics such as the Dice coefficient and 95% Hausdorff Distance were computed using the seg-metrics 1.2.7 Python library [38].
@@ Acknowledgements
Magnetic field strength 1.5 171 (100.0) 715 (73.0) 136 (46.7) 64 (100.0) 1086 (72.1)